# On the effectiveness of discrete representations in sparse mixture of experts

### v1

@@ 4.2 VQMoE solves Representation Collapse by Design
Same as the Jacobian matrix of SMoE, the Jacobian matrix of VQMoE consists two terms: (1) $J_{1}$ depends on input token and experts to the final output; (2) $\sum_{j=1}^{N+K+2}o_{j}\bm{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss. We can see that $N+K+2>>N$, it implies that VQMoE is better than SMoE to solve the representation collapse issue. In theory, we can choose the number of codes to be approximately $d-N-2$ with a hashing index to experts to address the issue. However, this involves a trade-off with the computational resources required to learn the codebook.
@@ 5 Experiment
We conduct experiments to explore the following hypotheses: (i) VQMoE provides an effective SMoE training algorithm for LLMs; (ii) VQMoE delivers a robust and efficient solution during the fine-tuning phase; and (iii) VQMoE outperforms other routing methods in vision domain.
@@ 5.1 Experimental Settings
To answer the three above hypotheses, we conduct experiments on Vision, Language, and Time-series tasks. For Pre-training language models, we examine two common tasks in the training and evaluation of large language models: character-level language modeling using the enwik8 and text8 datasets (Mahoney, 2011), and word-level language modeling with the WikiText-103 (Merity et al., 2016) and One Billion Word datasets (Chelba et al., 2014). For Parameter-efficient fine-tuning, we consider pre-trained base models on enwik8 and efficient Fine-tuning it on a downstream task. We choose the SST-2 (Socher et al., 2013), SST-5 (Socher et al., 2013), IMDB (Maas et al., 2011), and BANKING77 (Casanueva et al., 2020) datasets. For vision tasks, we employ the Vision Transformer model (Dosovitskiy et al., 2021) with the state-of-the-art routing method and our method to train and evaluate the image classification task. Our experiments encompass five widely recognized image classification datasets: Cifar10, Cifar100 (Krizhevsky, 2009), STL-10 (Coates et al., 2011), SVHN (Netzer et al., 2011), ImageNet-1K(Deng et al., 2009).
@@ 5.2 Pre-training Language Models
Training tasks  We explore two common tasks in the training and evaluation of LLMs. First, character-level language modeling on the enwik8 or text8 datasets (Mahoney, 2011), which are common datasets to evaluate the model’s pre-training capabilities. We also consider the word-level language modeling task on WikiText-103 (Merity et al., 2016) and One Billion Word dataset (Chelba et al., 2014), a much larger and more challenging dataset, to test the models scaling capabilities. For all datasets, we follow the default splits of training-validation-testing. Second, we consider Fine-tuning the models on downstream applications to investigate the models’ capabilities of adapting to different domains. To this end, we consider pre-trained medium models on enwik8 and Fine-tuning them on a downstream task. We choose the SST-2 (Socher et al., 2013), SST-5 (Socher et al., 2013), IMDB (Maas et al., 2011), and BANKING77 (Casanueva et al., 2020) datasets, which are common NLP tasks to evaluate pre-trained models. Following Chen et al. (2023a), we freeze the router and only optimize the experts’ parameter in this experiment.
Configuration Enwik8 (BPC) Text8 (BPC) WikiText-103 (PPL) lm1b (PPL) Architecture Algorithm Base Large Base Large Base Large Base Large Transformer VQMoE 1.48 1.41 1.47 1.40 38.74 31.98 59.48 49.30 SMoE 1.49 1.41 1.49 1.40 39.50 32.30 60.88 51.30 SMoE-Dropout 1.82 2.22 1.70 1.89 72.62 107.18 97.45 159.09 XMoE 1.51 1.42 1.49 1.42 39.56 32.65 61.17 51.84 StableMoE 1.49 1.42 1.49 1.41 39.45 32.34 60.72 50.74 Transformer-XL VQMoE 1.19 1.08 1.28 1.17 29.48 23.85 56.85 48.70 SMoE 1.20 1.09 1.29 1.18 30.16 24.02 58.00 48.71 SMoE-Dropout 1.56 2.24 1.56 1.86 58.37 40.02 93.17 68.65 XMoE 1.21 1.09 1.28 1.17 30.34 24.22 58.33 50.64 StableMoE 1.20 1.10 1.28 1.19 29.97 24.19 58.25 49.17 ##\## Params 20M 210M 20M 210M 20M 210M 20M 210M
Table 1: BPC on the enwik-8 and text8 test sets; and perplexity on the Wikitext-103 and One Billion Word test sets. Lower is better, best results are in bold.
Models. For the language tasks, we follow the same settings as in SMoE-Dropout (Chen et al., 2023a). We consider two decoder-only architectures: (i) the standard Transformer (Vaswani et al., 2017); and (ii) and Transformer-XL (Dai et al., 2019a) with the same number of parameters as Transformer. We evaluate our method versus the state of art Sparse Mixture of Expert Layers such as StableMoE (Dai et al., 2022) and XMoE (Chi et al., 2022). We consider two model configurations: (i) base: with four SMoE blocks and 20M parameters; (ii) large: with twelve SMoE layers and 210M parameters. We emphasize that we are not trying to achieve state-of-the-art results due to the limited resource constraints. Instead, we evaluate the small and large models on various datasets to demonstrate the scalability and efficacy of our algorithm. Lastly, we conduct extensive investigations using the tiny model to understand the algorithm behaviours and their robustness to different design choices. Lastly, unless otherwise stated, we implement them with $K=2$ in the experiments.
@@
Table 1 presents the evaluation metrics comparing VQMoE with state-of-the-art approaches. We also show the performance progression of the base model on the validation set. Notably, across all methods, the Transformer-XL architecture consistently outperforms the standard Transformer on all datasets. While advanced strategies like XMoE and StableMoE tend to surpass vanilla SMoE when model complexity is increased (from small to medium) or more data is introduced (moving from enwik8 to WikiText-103 or One Billion Word), these improvements are often inconsistent or marginal. In contrast, VQMoE consistently outperforms all competitors across benchmarks (keeping in mind that the BPC metric is log-scaled), architectures, and also converges more quickly. This highlights VQMoE’s effectiveness in learning an efficient routing policy for the language modeling pre-training task.
@@
Q3: When does VQMoE outperform router methods in terms of robustness? A3: The lower hidden size of FFN.
Compared to the routing methods, VQMoE achieves competitive performance which only requires 80% number of parameters. Figure  2(a) and Figure  2(b) demonstrate the robustness of our method on the Enwik8 and Text8 datasets, respectively.
@@ 5.3 Parameter-Efficient Fine-Tuning
Q4: What is the biggest advantage of SMoE, compared to the conventional SMoE? A4: Parameter-Efficient Fine-Tuning.
@@
Dataset SST-2 SST-5 IMDB BANKING77 SST-2 SST-5 IMDB BANKING77
VQMoE 5.6145 82.6 41.1 89.5 84.8 83.3 42.0 89.1 85.3
SMoE 7.7620 82.1 39.5 89.3 82.6 80.8 40.4 88.6 80.2
SMoE-Dropout 7.7620 81.3 39.6 88.9 77.9 81.8 40.0 89.1 77.3
XMoE 7.7620 82.4 39.9 89.0 83.1 81.3 40.3 88.7 82.7
StableMoE 7.7620 82.2 40.4 89.1 82.7 82.5 41.1 88.5 78.6
Table 2: Accuracy of the model after fine-tuned on various datasets. Higher is better, best results are in bold.
We see that the discrete representation that VQMoE learns at the Pretraning stage  5.2 might consist of rich knowledge. To test this hypothesis, we use only the discrete representation for downstream tasks, allowing VQMoE to save 28% of computational resources compared to SMoE. Table 2 reports the accuracy of the models fine-tuned on the test sets of various datasets. Overall, we observe that VQMoE demonstrates strong transfer learning capabilities by achieving the highest accuracy on all datasets. Notably, on the more challenging datasets of SST-5 and BANKING77, which have fewer training samples or more classes, we observe larger performance gains from VQMoE versus the remaining baselines (over $5\%$ improvements compared to the second-best method). This result shows that VQMoE can learn a discrete representation that is not only good for pre-training but also exhibits strong transfer capabilities to various downstream tasks.
@@ 5.4 Vision
To make our performance comparison informative and comprehensive, we consider two kinds of baselines that are fairly comparable to VQMoE: (1) Dense Model (Vision Transformer)  (Dosovitskiy et al., 2021); (2) SoftMoE  (Puigcerver et al., 2024) - the most advanced MoE in Vision domain. We perform two configurations for training the Mixture of Experts: (1) small - 10 million parameters (10M); (2) large - 110 million parameters (110M). The result at Table  3 shows that VQMoE outperforms both Vision Transformer Dense (Dosovitskiy et al., 2021), SoftMoE (Puigcerver et al., 2024), , and other routing methods such as  (Dai et al., 2022),  (Chi et al., 2022) on six out of eight tasks across four image classification datasets. We conduct our experiments three times on four datasets (CIFAR-10, CIFAR-100, STL-10, and SVHN) using different seeds, reporting the average results along with the standard deviation. For the large-scale dataset ImageNet-1K, we perform a single run due to resource constraints. The average performance of our method surpasses other baselines and is more stable, as indicated by the low standard deviation.
@@
Table 3: Accuracy of models evaluated on vision datasets. Higher is better, best results are in bold.
@@ 5.5 In-depth Analysis
Consistent Score. Figure  3(a) illustrates that expert selections when training SMoE face inconsistent problems. As the Theorem  4.1, this inconsistency arises because the router’s coverage rate significantly exceeds that of the Transformer representation. The figure  3(a) also shows that our method achieves the highest consistency score compared to the SMoE and XMoE models. However, the VQMoE model’s consistency score is around 75%, as our method also requires learning a continuous representation during the Pre-training phase.
Representation Collapse issue. To visualize the Representation collapse problem in practice, we apply Principal Component Analysis (PCA) method to reduce from $d$ dimension of the Transformer to 2D for plotting purposes, thanks to  (Chi et al., 2022). Figures  3(b) and  3(c) show the expert representations from the pretrained VQMoE and SMoE models. The results suggest that VQMoE experiences less representation collapse in the expert space compared to SMoE. The analysis is in line with the theorem proof at Section  4.2. However, projecting the $d$-dimensional space onto 2D for visualization may lead to information loss.
@@
Figure 3: Analysis Inconsistent Expert Selection and Representation Collapse issues when training SMoE. Figure  3(a) demonstrates consistent score movement from VQMoE, compared with SMoE and XMoE. Figure  3(b) and Figure  3(c) visualize the representation by experts in 2D dimension using Principal Component Analysis (PCA) method.
@@ 5.6 Ablation Study
Sensitiveness of VQ loss contribution $\alpha$. Figure 6(c) illustrates the impact of $\alpha$, which controls the contribution of the Vector Quantization loss to the overall loss. If $\alpha$ is too high, it leads to a better discrete representation but may negatively affect the final target. Conversely, if $\alpha$ is too low, it may result in a poor discrete representation. Therefore, $\alpha$ should be selected based on the data, typically within the range of $(0.05,0.15)$.
@@ 6 Conclusion and Future Directions
This study illustrates Vector-Quantized Mixture of Experts (VQMoE), which is novel and theoretically-grounded architecture to overcome challenges in training SMoE such as representation collapse and inconsistency. We evaluate our method on various Pre-training and Fine-tuning tasks, for both language and vision domains. The results show that VQMoE outperforms the routing methods both theoretically and empirically. Furthermore, fine-tuning VQMoE with the discrete representation for downstream tasks could reduce computational resource usage by 28%. We believe that focusing on discrete representation learning will offer a promising strategy for training and testing sparse mixtures of experts (SMoE) at a large scale. Finally, we believe that our approach opens up new research avenues for effectively training SMoE, where cutting-edge techniques in discrete representation learning and vector quantization can be harnessed to enhance their performance.
@@ Appendix A Appendix
Supplementary Material for “On the effectiveness of discrete representations in sparse mixture of experts"
This document is organized as follow. Appendix B presents the detailed proof of our theoretical analysis in Section 4. Appendix C provide in-depth analysis about the representation collapse while Appendix D presents all the implementation details and additional results.
Appendix B Proof for Results in Section 4
B.1 Proof of Theorem 4.1
Definition B.1 (Consistent Router)
A sequence of points $x_{1},x_{2},\ldots,x_{n}$ and a corresponding sequence of clusters $C_{1},C_{2},\ldots,C_{k}$ are said to be consistent if, for every point $x_{p}\in C_{i}$, the condition
$\text{dist}(x_{p},u_{i})\leq\min_{j\neq i}\text{dist}(x_{p},u_{j})$
Definition B.2 (Inconsistent Router)
A sequence of points $x_{1},x_{2},\ldots,x_{n}$ and a corresponding sequence of clusters $C_{1},C_{2},\ldots,C_{k}$ are said to be inconsistent if there exists a point $x_{p}\in C_{i}$ such that
$\text{dist}(x_{p},u_{i})>\min_{j\neq i}\text{dist}(x_{p},u_{j}),$
In this proof, we use contradiction to establish the theorem. Assume that the expert embeddings $e$ form a consistent router. By Definition B.1, we have:
$\text{dist}(x_{p},u_{i})\leq\min(\text{dist}(x_{p},C_{j})),$

### v2

@@ 4.2 Experts Representation Collapse
Same as the Jacobian matrix of SMoE, the Jacobian matrix of VQMoE consists two terms: (1) $J_{1}$ depends on input token and experts to the final output; (2) $\sum_{j=1}^{N+K+2}o_{j}\bm{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss. We can see that $N+K+2>>N$, implying that VQMoE is better than SMoE in solving the representation collapse issue. In theory, we can choose the number of codebook to be approximately $d-N-2$ with a hashing index to experts to address the issue. However, this involves a trade-off with the computational resources required to learn the codebook.
@@ 5 Experiment
We conduct experiments to investigate the following hypotheses: (i) VQMoE offers an effective training algorithm for Sparse Mixture-of-Experts (SMoE) in large language models (LLMs); (ii) VQMoE enables efficient fine-tuning; and (iii) VQMoE outperforms other routing methods across multiple domains.
@@ 5.1 Experimental Settings
To evaluate the three hypotheses, we conduct experiments across both vision and language tasks. For pre-training language models, we assess two standard benchmarks: (i) character-level language modeling using enwik8 and text8(Mahoney, 2011), and (ii) word-level language modeling using WikiText-103(Merity et al., 2016) and the more challenging One Billion Word (lm1b) dataset (Chelba et al., 2014). All experiments use the standard training, validation, and test split with a 90:5:5 ratio as(Child et al., 2019).
For parameter-efficient fine-tuning, we fine-tune models pre-trained on enwik8 using four widely used NLP datasets: SST-2, SST-5(Socher et al., 2013), IMDB(Maas et al., 2011), and BANKING77(Casanueva et al., 2020). FollowingChen et al. (2023a), we freeze the router and update only the expert parameters to evaluate fine-tuning efficiency.
For vision tasks, we employ the Vision Transformer (ViT)(Dosovitskiy et al., 2021) and compare our routing method with state-of-the-art alternatives on five benchmark image classification datasets: CIFAR-10, CIFAR-100(Krizhevsky, 2009), STL-10(Coates et al., 2011), SVHN(Netzer et al., 2011), and ImageNet-1K (Deng et al., 2009).
@@ 5.2 Pre-training Language Models
Configuration Enwik8 (BPC) Text8 (BPC) WikiText-103 (PPL) lm1b (PPL) Avg. Char-level Avg.Word-level Architecture Algorithm Base Large Base Large Base Large Base Large - - Transformer VQMoE 1.48 1.41 1.47 1.40 38.74 31.98 59.48 49.30 1.44 44.88 SMoE 1.49 1.41 1.49 1.40 39.50 32.30 60.88 51.30 1.45 45.50 SMoE-Dropout 1.82 2.22 1.70 1.89 72.62 107.18 97.45 159.09 1.91 109.59 XMoE 1.51 1.42 1.49 1.42 39.56 32.65 61.17 51.84 1.46 46.06 StableMoE 1.49 1.42 1.49 1.41 39.45 32.34 60.72 50.74 1.45 45.81 Transformer-XL VQMoE 1.19 1.08 1.28 1.17 29.48 23.85 56.85 48.70 1.18 39.72 SMoE 1.20 1.09 1.29 1.18 30.16 24.02 58.00 48.71 1.19 40.22 SMoE-Dropout 1.56 2.24 1.56 1.86 58.37 40.02 93.17 68.65 1.81 65.55 XMoE 1.21 1.09 1.28 1.17 30.34 24.22 58.33 50.64 1.19 40.88 StableMoE 1.20 1.10 1.28 1.19 29.97 24.19 58.25 49.17 1.19 40.40 #\## Params 20M 210M 20M 210M 20M 210M 20M 210M - -
Table 1: Bits-per-character (BPC) on the Enwik8 and Text8 test sets; and perplexity (PPL) on the WikiText-103 and One Billion Word test sets. Avg. Char-level is the average BPC over Enwik8 and Text8; Avg. Word-level is the average PPL over WikiText-103 and lm1b. Lower is better; best results are in bold.
[…]
@@
(a) Training PPL movement on Wikitext-103 dataset.
(b) Training PPL movement on lm1b dataset.
Figure 2: Perplexity (PPL) over training steps for the Transformer-XL base model on two datasets: (a) WikiText-103 and (b) lm1b. The results indicate that VQMoE converges faster than the baseline models, demonstrating its efficiency and robustness for language modeling tasks.
@@
Table 1 presents the evaluation metrics comparing VQMoE with state-of-the-art approaches. We also show the performance progression of the base model on the validation set. Notably, across all methods and datasets, VQMoE consistently outperforms the baseline models for both the Transformer-XL and Transformer architectures on average. Although advanced strategies such as XMoE and StableMoE generally outperform the vanilla SMoE on character-based datasets such as enwik8 and text8, which involve a small vocabulary size, their improvements tend to diminish or become marginal when trained on more complex, large-vocabulary datasets such as WikiText-103 and One Billion Word (lm1b). In contrast, VQMoE consistently outperforms all competitors across benchmarks (keeping in mind that the BPC metric is log-scaled), architectures, and also converges more quickly as Figure 2. This highlights VQMoE’s effectiveness in learning an efficient routing policy for the language modeling pre-training task.
@@
Q3: Can VQMoE, with only 80% of the total parameter count, achieve better performance than SMoE utilizing the full 100% of parameters? A3: Yes.
To evaluate the robustness of VQMoE, we reduce its hidden dimension to half that of the SMoE baseline, resulting in approximately a 20% reduction in the total number of parameters. Robustness here denotes the model’s ability to maintain strong performance across different parameter scales, particularly with fewer parameters. We then train both models across a range of parameter scales: 1M, 2M, 4M, 8M, and 16M, where M denotes the number of parameters in millions. Despite having only 80% of the parameter count, VQMoE consistently achieves competitive performance compared to SMoE across all scales. This highlights the efficiency and robustness of our approach. The results are illustrated in Figure 3(a) and Figure 3(b), which show VQMoE’s performance on the Enwik8 and Text8 datasets, respectively.
@@ 5.3 Parameter-Efficient Fine-Tuning
Q4: What is the biggest advantage of VQMoE, compared to the conventional SMoE? A4: Parameter-Efficient Fine-Tuning.
@@
Dataset SST-2 SST-5 IMDB BANKING77 SST-2 SST-5 IMDB BANKING77 -
VQMoE 5.6145 82.6 41.1 89.5 84.8 83.3 42.0 89.1 85.3 74.72
SMoE 7.7620 82.1 39.5 89.3 82.6 80.8 40.4 88.6 80.2 72.94
SMoE-Dropout 7.7620 81.3 39.6 88.9 77.9 81.8 40.0 89.1 77.3 72.00
XMoE 7.7620 82.4 39.9 89.0 83.1 81.3 40.3 88.7 82.7 73.43
StableMoE 7.7620 82.2 40.4 89.1 82.7 82.5 41.1 88.5 78.6 73.89
Table 2: Accuracy of the model after fine-tuning on various datasets. Higher is better; best results are in bold.
[…]
@@ 5.4 Vision
To make our performance comparison informative and comprehensive, we consider two kinds of baselines that are fairly comparable to VQMoE: (1) Dense Model (Vision Transformer)  (Dosovitskiy et al., 2021); (2) SoftMoE  (Puigcerver et al., 2024) - the most advanced MoE in Vision domain. We perform two configurations for training the Mixture of Experts: (1) small - 10 million parameters (10M); (2) large - 110 million parameters (110M). The result at Table  3 shows that VQMoE outperforms both Vision Transformer Dense (Dosovitskiy et al., 2021), SoftMoE (Puigcerver et al., 2024), and other routing methods such as  (Dai et al., 2022),  (Chi et al., 2022) on six out of eight tasks across four image classification datasets. We conduct our experiments three times on four datasets (CIFAR-10, CIFAR-100, STL-10, and SVHN) using different seeds, reporting the average results along with the standard deviation. For the large-scale dataset ImageNet-1K, we perform a single run due to resource constraints. The average performance of our method surpasses other baselines and is more stable, as indicated by the low standard deviation.
@@
Table 3: Accuracy of models evaluated on vision datasets. Higher is better, the best results are in bold.
@@ 5.5 In-depth Analysis
Consistent Score. Figure  4(a) illustrates that expert selections when training SMoE face inconsistent problems. As the Theorem  4.3, this inconsistency arises because the router’s coverage rate significantly exceeds that of the Transformer representation. Figure  4(a) also shows that our method achieves the highest consistency score compared to the SMoE and XMoE models. However, the VQMoE model’s consistency score is around 75%, as our method also requires learning a continuous representation during the Pre-training phase.
Representation Collapse issue. To visualize the Representation collapse problem in practice, we apply Principal Component Analysis (PCA) method to reduce from $d$ dimension of the Transformer to 2D for plotting purposes, thanks to  (Chi et al., 2022). Figures  4(b) and  4(c) show the expert representations from the pretrained VQMoE and SMoE models. The results suggest that VQMoE experiences less representation collapse in the expert space compared to SMoE. The analysis is in line with the theorem proof at Section  4. However, projecting the $d$-dimensional space onto 2D for visualization may lead to information loss.
@@
Figure 4: Analysis Inconsistent Expert Selection and Representation Collapse issues when training SMoE. Figure  4(a) demonstrates consistent score movement from VQMoE, compared with SMoE and XMoE. Figure  4(b) and Figure  4(c) visualize the representation by experts in 2D dimension using Principal Component Analysis (PCA) method.
@@ 5.6 Ablation Study
Sensitiveness of VQ loss contribution $\alpha$. Figure 6(c) illustrates the impact of $\alpha$, which controls the contribution of the Vector Quantization loss to the overall loss. If $\alpha$ is too high, it leads to a better discrete representation but may negatively affect the final target. Conversely, if $\alpha$ is too low, it may result in a poor discrete representation. Therefore, $\alpha$ should be selected based on the data, typically within the range of (0.05,0.15)(0.05,0.15)( 0.05 , 0.15 ).
@@ 6 Conclusion and Future Directions
This study illustrates Vector-Quantized Mixture of Experts (VQMoE), a novel and theoretically-grounded architecture, to overcome challenges in training SMoE such as representation collapse and inconsistency. We evaluate our method on various Pre-training and Fine-tuning tasks, for both language and vision domains. The results show that VQMoE outperforms the routing methods both theoretically and empirically. Furthermore, fine-tuning VQMoE with the discrete representation for downstream tasks could reduce computational resource usage by 28%. We believe that focusing on discrete representation learning will offer a promising strategy for training and testing sparse mixtures of experts (SMoE) at a large scale. Finally, we believe that our approach opens up new research avenues for effectively training SMoE, where cutting-edge techniques in discrete representation learning and vector quantization can be harnessed to enhance their performance.
@@ Appendix A Appendix
Supplementary Material for “On the Role of Discrete Representation in Sparse Mixture of Experts”
This document is organized as follows. Appendix A.1 provides a detailed proof for Section 4. Appendix A.2 presents additional experimental results demonstrating the effectiveness of our method compared to the baselines. Finally, Appendix A.3 offers an in-depth analysis of representation collapse, while Appendix A.4 details the implementation aspects.
A.1 Proof for Results in Section 4
A.1.1 Jacobian Matrix of VQMoE
To investigate whether VQMoE alleviates this collapse, we derive the Jacobian of the VQMoE output with respect to the input $x\in\mathbb{R}^{n\times d}$:
$\bm{J}^{\text{VQMoE}}$ $=g_{c}(\bm{x})\cdot\bm{J}^{\text{SMoE}}+\frac{\partial g_{c}(\bm{x})}{\partial\bm{x}}f^{\mathrm{SMoE}}(\bm{x})$ (12)
$\quad+g_{d}(\bm{x})\cdot\sum_{l=1}^{K}\bm{J}^{\text{FFN}}_{l}+\frac{\partial g_{d}(\bm{x})}{\partial\bm{x}}\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\tilde{\bm{x}}_{l})$
$=g_{c}(\bm{x})\cdot\left[J_{1}+\sum_{j=1}^{N}\bm{c}_{j}\bm{e}_{j}^{\top}\right]+\sum_{m\in\{c,d\}}g_{m}\bm{e}_{m}^{\top}+\sum_{l=1}^{K}d_{l}\bm{e}_{l}^{\top}$
$=J_{1}+\sum_{j=1}^{N}c_{j}\bm{e}_{j}^{\top}+\sum_{l=1}^{K}d_{l}\bm{e}_{l}^{\top}+\sum_{m\in\{c,d\}}g_{m}\bm{e}_{m}^{\top}$
$=J_{1}+\sum_{j=1}^{N+K+2}o_{j}\bm{e}_{j}^{\top}.$
where:
[…]
As in SMoE, the Jacobian of VQMoE consists of two major components: $J_{1}$ the primary contribution from the input and selected expert; and $\sum_{i=1}^{N+K+2}o_{i}\bm{e}_{i}^{\top}$ additional gradient contributions from both the continuous part and the discrete part.
A.1.2 Proof of Theorem 4.3
In this proof, we use contradiction to establish the theorem. Assume that the expert embeddings $e$ form a consistent router. By Definition 4.1, we have:
$\text{dist}(x_{p},u_{i})\leq\min(\text{dist}(x_{p},u_{j})),$